A private shipping company will accept a box for domestic shipment only if the sum of its length and girth (distance around) does not exceed 108 in. Suppose you want to mail a box with square sides so that its dimensions are   and it's girth is   What dimensions will give the box its largest volume?




1
Expert's answer
2021-03-01T06:42:39-0500

First, we will need to complete the question.

Let us assume that its dimensions are h by h by w and its girth is 2h + 2w. 

"Volume=h^2w"

Length + girth =108

"w+(2h+2w)=108"

"2h+3w=108 \implies w = \frac{108-2h}{3} \implies w = 36-\frac{2h}{3}"

"V=h^2w=h^2(36-\frac{2h}{3})=36h^2-\frac{2h^3}{3}"

Maximum volume "\frac{dV}{dh}=0 \implies\frac{dV}{dh}=0=72h-2h^2 \implies h=36"

"w=36−\frac{2}{3}*36=12"

The dimension that gives the box its largest volume is "36\times36\times12"